OpenStudy (anonymous):

How do I solve this with quadratic formula 2x(x-1)=3

JoelTheBoss (joel_the_boss):

@bellablain

OpenStudy (bellablain):

I assume you want to solve for x 2(x-1) + 3 = x - 3(x+1) multiply the 2 and -3 into the (x - 1) and (x + 3) 2x -2 + 3 = x - 3x - 3 combine like terms 2x + 1 = -2x - 3 add 2x to both side, subtract 1 from both sides 4x = -4 divide both sides by 4 x = -1 verify by putting -1 in for x in the orig equation 2(x-1) + 3 = x - 3(x+1) 2(-1 - 1) + 3 = -1 - 3(-1 + 1) 2(-2) + 3 = -1 - 3(0) -4 + 3 = -1 -1 = -1

OpenStudy (anonymous):

I have to make them equal to 0

OpenStudy (bellablain):

um i hope this helps i dont know what you need. Simplifying 2x(x + -1) = 3 Reorder the terms: 2x(-1 + x) = 3 (-1 * 2x + x * 2x) = 3 (-2x + 2x2) = 3 Solving -2x + 2x2 = 3 Solving for variable 'x'. Reorder the terms: -3 + -2x + 2x2 = 3 + -3 Combine like terms: 3 + -3 = 0 -3 + -2x + 2x2 = 0 Begin completing the square. Divide all terms by 2 the coefficient of the squared term: Divide each side by '2'. -1.5 + -1x + x2 = 0 Move the constant term to the right: Add '1.5' to each side of the equation. -1.5 + -1x + 1.5 + x2 = 0 + 1.5 Reorder the terms: -1.5 + 1.5 + -1x + x2 = 0 + 1.5 Combine like terms: -1.5 + 1.5 = 0.0 0.0 + -1x + x2 = 0 + 1.5 -1x + x2 = 0 + 1.5 Combine like terms: 0 + 1.5 = 1.5 -1x + x2 = 1.5 The x term is -1x. Take half its coefficient (-0.5). Square it (0.25) and add it to both sides. Add '0.25' to each side of the equation. -1x + 0.25 + x2 = 1.5 + 0.25 Reorder the terms: 0.25 + -1x + x2 = 1.5 + 0.25 Combine like terms: 1.5 + 0.25 = 1.75 0.25 + -1x + x2 = 1.75 Factor a perfect square on the left side: (x + -0.5)(x + -0.5) = 1.75 Calculate the square root of the right side: 1.322875656 Break this problem into two subproblems by setting (x + -0.5) equal to 1.322875656 and -1.322875656. Subproblem 1 x + -0.5 = 1.322875656 Simplifying x + -0.5 = 1.322875656 Reorder the terms: -0.5 + x = 1.322875656 Solving -0.5 + x = 1.322875656 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '0.5' to each side of the equation. -0.5 + 0.5 + x = 1.322875656 + 0.5 Combine like terms: -0.5 + 0.5 = 0.0 0.0 + x = 1.322875656 + 0.5 x = 1.322875656 + 0.5 Combine like terms: 1.322875656 + 0.5 = 1.822875656 x = 1.822875656 Simplifying x = 1.822875656 Subproblem 2 x + -0.5 = -1.322875656 Simplifying x + -0.5 = -1.322875656 Reorder the terms: -0.5 + x = -1.322875656 Solving -0.5 + x = -1.322875656 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '0.5' to each side of the equation. -0.5 + 0.5 + x = -1.322875656 + 0.5 Combine like terms: -0.5 + 0.5 = 0.0 0.0 + x = -1.322875656 + 0.5 x = -1.322875656 + 0.5 Combine like terms: -1.322875656 + 0.5 = -0.822875656 x = -0.822875656 Simplifying x = -0.822875656 Solution The solution to the problem is based on the solutions from the subproblems. x = {1.822875656, -0.822875656}

OpenStudy (anonymous):

Thanks but I'm looking for the quadratic formula

OpenStudy (bellablain):